# Maxime Dieudonné1, Guillaume Auzias1,†, Julien Lefèvre1,†

### v1

@@ 3.2 Experiment #1 : setting of the parameter alpha of the $DPF^{*}$
The aim of this experiment is to assess the influence of the $\alpha$ parameter of $DPF^{*}$ and to define its most efficient setting. To do so, we computed the DPF* with various values of $\alpha$ ranging from 0 to 2000 on the set of selected brains described below. We also computed the SULC and DPF estimations, as well as the mean curvature (denoted as CURV) using the method described in [Rusinkiewicz,] in order to quantitatively compare the methods on the same set of surfaces.
@@
    where $std(D\rvert_{\mathcal{C}})$ is the standard deviation of $D$ across vertices located on the gyral crests. Since we want to evaluate the variation in depth values on the crests in relation to the variation in depth across all the vertices of $M$ (and not only on $\mathcal{C}$), we normalise according to $IPR(D)$. $IPR(.)$ stands for inter-percentile range with $percentile(D,x)$ being the $x$th percentile of the depth distribution across all the vertices of $M$. $IPR(.)$ is a robust estimation of the spread of a distribution [Ross,]. Using this metric, an ideal sulcal depth map would have a $StdCrest$ value close to 0.
@@
    $Sep(D)=\frac{1}{IPR(D)}(median(D\rvert_{\mathcal{C}})-median(D\rvert_{\mathcal{F}})$ (22)
    $Sep$ varies theoretically between 0 and 1. A value of 1 corresponds to the ideal case where the crests have a constant and maximal depth value and the fundi a minimal and constant value. This is not realistic as the fundi are not expected to have a constant depth value. We also expect that the range of values on the fundi is larger for adult brains compared to less folded baby brains due to the appearance of secondary and tertiary sulci. In any case, a low $Sep$ (below 0.5) would be indicative of inconsistent sulcal depth values, with relatively high values even in the fundi and/or relatively low values on the gyral crests.
@@
We report in Fig.2 and corresponding Sec4.1 the distribution of these three measures across all subjects as violin plots. Independently for each measure, we then identified the best value for $\alpha$ based on the median across all subjects (lowest median for $StdCrest$ and $Dev$, highest for $Sep$). For each measure, we then performed a t-test between each distribution obtained for the different values of $\alpha$ and the distribution corresponding to the best $\alpha$, allowing us to define a range of values on which the measures are statistically equivalent. Finally, we identify the optimal value for $\alpha$ across the three measures by looking for values that fall in the intersection between the three ranges of values that are equivalent to the best one. This optimal value for $\alpha$ is then used in the other experiments.
@@ 3.4 Experiment #3 : Application to a large dataset
We computed SULC, the $DPF^{*}$ and $DPF^{*}_{abs}$ for each of these surfaces and show the distributions across the vertices of the corresponding mesh of the resulting sulcal depth measures as a function of the surface area. In addition, we computed the Wasserstein distance as proposed in Eq.8, between every pair of individuals, resulting in a matrix of Wasserstein distances for each method. As explained in Sec.2.1, this allows us to quantify the deviation to the scale controlled case for the $DPF^{*}$, and to quantify the influence of global brain size on the estimated sulcal depth measures. We then assessed whether the Wasserstein distances computed for larger brains were statistically different from the distances computed for smaller brains. To do so, we selected 8 subgroups of 200 subjects corresponding to subpopulations of different brain size, ordered from smaller brains to larger brains. This approach corresponds to a sliding window analysis of the patterns observed in the distance matrices. We then set as a reference subgroup the 200 subjects with larger brain size and assessed statistically putative differences in the distribution of Wasserstein distances between this subpopulation and the other subgroups using the Kolmogorov-Smirnov statistic for two samples. The resulting statistic, with values between 00 and $1$, indicates how similar are the distributions from the other subgroups with respect to the group with larger brains. In other words this statistic is a quantitative measure of the influence of the variations in brain size on the estimated sulcal depth maps: A low statistic value would mean that the sulcal depth maps estimated for the two subgroups are similar despite variations in global brain size. A high statistic value would correspond to the case where the distributions are very different.
@@ 4.1 Result from experiment #1: setting of parameter $\alpha$
The results of Expe #1 are shown in Fig.2 On this figure, we report the distributions from the three measures detailed in Sec.3.2 ($Dev$, $StdCrest$, and $Sep$) computed across the dataset #1, for different values of the $\alpha$ parameter.
Figure 2: Results from Expe #1. A: From top to bottom, we show the distributions of the three measures $Dev$, $StdCrest$, and $Sep$ across all the brains from the dataset #1 as violin plots, for the different values of $\alpha$. We highlight in blue the range of values for which the distributions are statistically equivalent to the one obtained using the best value for $\alpha$ (t-test with p-value$>$.05). B: For each measure, we show the distributions obtained for the method SULC, the mean curvature and the classical $DPF$ from [Auzias et al., a], for compmarison.
@@
From the plot of the normalised standard deviation on the crests $StdCrest$, we observe a strong decrease with increasing $\alpha$, from .24 for $\alpha=0$ to .12 for $\alpha=150$. The lowest median is obtained for $\alpha=400$, and the statistical tests show that the distributions are equivalent for $\alpha\in[150,1000]$.
From the plot of the difference between the median of the crest and the median of the fundi $Sep$, we observe a progressive increase with increasing values of $\alpha$. From a value of .2 for $\alpha=0$, the median of $Sep$ increases up to .9 for $\alpha=500$, with a highest median for $\alpha=2000$. The statistical tests indicate that the distributions are equivalent for $\alpha$ values above 500.
@@
When comparing with the SULC and classical $DPF$ methods in Fig.1.B, we observe that their angular deviation $Dev$ are also quite low compared to mean curvature, and not statistically different from the best setting of $\alpha$ for $DPF^{*}$. These distributions confirm the visual observation of the appropriate setting of the default parameters of SULC for brains having a global scale not too different from an adult. Note however that for $StdCrest$ and $Sep$, the distributions from $DPF^{*}$ with the best $\alpha$ are statistically different from SULC, and from the classical $DPF$.
Overall, we observe that the influence of $\alpha$ on the estimated suclal depth from $DPF^{*}$ is continuous, since a small variation of $\alpha$ will have a very small influence. In addition, $DPF^{*}$ shows better measures than both DPF and SULC for a large range of values of $\alpha$. When considering jointly the three measures, we identify a quasi-optimal setting for $\alpha=500$.
@@ 4.2 Result from experiment #2: invariance to scaling
Figure 3: Results from Expe #2. A. The scaled surfaces obtained as described in Sec.3.3. B. Depth maps generated using the $DPF^{*}$ method (with $\alpha=500$). The scale-invariance is confirmed empirically: the estimated depth values are not influenced by the scaling. The correlation is thus perfect for all surfaces and the regression lines are super-imposed. C. Left: Depth estimations resulting from the SULC method applied to the scaled surfaces. Middle: both the slope and the correlation values depend on the scaling factor. Right: Spatial distribution of residuals from the linear regression models. Consistently with the regression plot, the residuals increase with the scaling factor. The regions corresponding to most concave and convex geometric configurations show higher magnitude of error.
First, this experiment allows us to confirm empirically that the $DPF^{*}$ is scale-invariant: $D^{*}(M,s)=D^{*}(M,1)$ for $s={2,3,4,5}$. This experimental confirmation aligns with the theoretical formulation we introduced earlier.
@@ 4.3 Result from experiment #3: Application to a large dataset
The results of Expe #3 are detailed in Fig.4 and corresponding caption. The distributions of the different sulcal depth estimation techniques (top row) are indicative of the influence of the global brain size. For SULC, on the right, the impact of the increase in brain size from left to right is visible on all the centiles. More specifically, the jump in the dispersion of the distributions (esp. visible on the 75th and 95th centiles) corresponds to the transition from the 879 newborns from the dHCP to the adults from the HCP. On the dHCP dataset, the dispersion progressively increases with brain size, while on the HCP dataset, the distributions are very similar across individuals despite variations in global brain size (52% of variation between the volumes of the smallest and largest adults from HCP).
@@
For the $DPF^{*}_{abs}$, at the centre, we see that by multiplying the $DPF^{*}$ by the characteristic length as in Eq.14, we obtain an absolute estimation of the sulcal depth with a clear influence of the global size of the mesh, as expected. The increase in the dispersion of the distributions with brain size is clearly visible and progressive. Of note, the influence of brain size is stronger than for SULC on the adult subjects from the HCP dataset. The $DPF^{*}_{abs}$ shows a higher sensitivity than SULC to the 52% variation in brain volume across adults.
@@
Figure 4: Analysis of the distribution of $DPF^{*}$, $DPF^{*}_{abs}$ and SULC across the 879 surfaces from the dHCP newborns and 1108 surfaces from HCP young adults. From left to right, the three columns correspond resp. to $DPF^{*}$, $DPF^{*}_{abs}$ and SULC. On the top row, we show the distribution of each sulcal depth estimation for all subjects, ordered following their global size. For each method, we show the mean in purple, the median in blue, the 25th and 75th centiles in green, and the 5th and 95th centiles in orange. On the second row, we show the corresponding matrices of the Wasserstein distance between every pair of surfaces. On the third row, we show for each method the distributions of the Wasserstein distances within the four sub-populations corresponding to the blocks illustrated on the matrix for the method SULC. At the bottom, we show for each method the KS statistic comparing the distribution in the block corresponding to the subgroup of largest brains, to the other blocks corresponding to smaller brains. See the text for further description.

### v2

@@ 3.1 Experiment #1 : setting of the parameter alpha of the $DPF^{*}$
The aim of this experiment is to assess the influence of the $\alpha$ parameter of $DPF^{*}$ and to define its most efficient setting. To do so, we computed the DPF* with various values of $\alpha$ ranging from 0 to 2000 on the set of selected brains described below. We also computed the SULC and DPF estimations, as well as the mean curvature (denoted as CURV) using the method described in rusinkiewicz_estimating_2004 in order to quantitatively compare the methods on the same set of surfaces.
@@
    where $std(D\rvert_{\mathcal{C}})$ is the standard deviation of $D$ across vertices located on the gyral crests. Since we want to evaluate the variation in depth values on the crests in relation to the variation in depth across all the vertices of $M$ (and not only on $\mathcal{C}$), we normalise according to $IPR(D)$. $IPR(.)$ stands for inter-percentile range with $percentile(D,x)$ being the $x$th percentile of the depth distribution across all the vertices of $M$. $IPR(.)$ is a robust estimation of the spread of a distribution ross_introductory_2017. Using this metric, an ideal sulcal depth map would have a $StdCrest$ value close to 0.
  • 2.
@@
[…]
    SepSep varies theoretically between 0 and 1. A value of 1 corresponds to the ideal case where the crests have a constant and maximal depth value and the fundi a minimal and constant value. This is not realistic as the fundi are not expected to have a constant depth value. We also expect that the range of values on the fundi is larger for adult brains compared to less folded baby brains due to the appearance of secondary and tertiary sulci. In any case, a low $Sep$ (below 0.5) would be indicative of inconsistent sulcal depth values, with relatively high values even in the fundi and/or relatively low values on the gyral crests.
  • 3.
@@
We report in Fig.2 and corresponding Sec4.1 the distribution of these three measures across all subjects as violin plots. Independently for each measure, we then identified the best value for $\alpha$ based on the median across all subjects (lowest median for $StdCrest$ and $Dev$, highest for $Sep$). For each measure, we then performed a non-parametric Wilcoxon signed-rank test between each distribution obtained for the different values of $\alpha$ and the distribution corresponding to the best $\alpha$, allowing us to define a range of values on which the measures are statistically equivalent.
@@ 3.3 Experiment #3 : Application to a large dataset
We computed SULC, the $DPF^{*}$ and $DPF^{*}_{abs}$ for each of these surfaces and show the distributions across the vertices of the corresponding mesh of the resulting sulcal depth measures as a function of the surface area. In addition, we computed the Wasserstein distance as proposed in Eq.8, between every pair of individuals, resulting in a matrix of Wasserstein distances for each method. As explained in Sec.2.1, this allows us to quantify the deviation to the scale controlled case for the $DPF^{*}$, and to quantify the influence of global brain size on the estimated sulcal depth measures. We then assessed whether the Wasserstein distances computed for larger brains were statistically different from the distances computed for smaller brains. To do so, we selected 8 subgroups of 200 subjects corresponding to subpopulations of different brain size, ordered from smaller brains to larger brains. This approach corresponds to a sliding window analysis of the patterns observed in the distance matrices. We then set as a reference subgroup the 200 subjects with larger brain size and assessed statistically putative differences in the distribution of Wasserstein distances between this subpopulation and the other subgroups using the Kolmogorov-Smirnov statistic for two samples. The resulting statistic, with values between 0 and $1$, indicates how similar are the distributions from the other subgroups with respect to the group with larger brains. In other words this statistic is a quantitative measure of the influence of the variations in brain size on the estimated sulcal depth maps: A low statistic value would mean that the sulcal depth maps estimated for the two subgroups are similar despite variations in global brain size. A high statistic value would correspond to the case where the distributions are very different.
3.4 Experiment #4 : Quantifying brain folding dynamics
In this experiment, we compare SULC and $DPF^{*}_{abs}$ at a local scale by assessing the dynamics of the formation of the main cortical suci. To do so, we applied the different estimation techniques to the spatio-temporal surfacic atlas computed from the dHCP dataset as described in bozek_construction_2018, and compared the depth values in 23 sulcal regions of interest distributed across the entire cortex. This atlas is representative of the evolution of the cortical surface between 28 and 44 weeks post-conception and thus shows marked changes in both the global size and in the magnitude of folding. For each of the 16 time steps available in the atlas, we applied SULC and $DPF^{*}_{abs}$ independently. The delineation of the sulcal regions was done by semi-automated tracing on the oldest atlas (44 weeks) as follows: We first extracted all the sulcal basins defined as the regions with negative values in the smoothed mean curvature map. We then manually labeled a set of sulcal basins according to the nomenclature provided in de_vareilles_development_2023, aiming at covering most of the cortex, with sulci expected to appear at different gestational ages. We did not label the sulcal basins corresponding to very small sulci for which the correspondence across the different time steps of the atlas would not be anatomically accurate. We obtained 23 sulcal regions corresponding to sulci appearing across the entire age range covered by the atlas. Since the atlas was designed using surface-based registration ensuring correspondence in both space and time (a given vertex on the mesh can be mapped across the different time steps, see bozek_construction_2018 for details), the sulcal regions defined in the last time step (44 weeks) can be projected onto earlier time steps, and the anatomical relevance of the definition of the sulcal regions can be assessed visually, as shown on Fig.5. This approach allowed us to delineate the 23 sulcal regions for all the time steps of the atlas, and we were able to track the evolution with age of the estimated depth from the different techniques.
3.5 Experiment #5 : Assessment of the potential influence of the spatial resolution of the mesh
In this experiment, we assess the potential impact on the $DPF^{*}$ of variations in the spatial resolution of the mesh, i.e. in the number of vertices and triangles in the mesh. To do so, we use the 16 cortical surfaces of experiment #1 that cover a large range of levels of cortical folding, and apply a the down sampling scheme garland1997surface as implemented in the PyMeshlab package pymeshlab. For each surface we obtain 3 levels of lower resolution by keeping 25, 50, 75 % of the vertices. We then compute the $DPF^{*}$ for all the surfaces and project the depth map computed on down-sampled meshes onto the corresponding original, high-resolution, mesh by using a nearest neighbor interpolation in 3D. We finally compute the normalized root mean squared error (NRMSE) expressed as a percentage of the value from the original mesh:
$NRMSE(DPF_{orig},DPF_{proj})=100\times\frac{\sqrt{\frac{1}{N_{low}}\sum_{i=1}^{N_{low}}(DPF_{orig}(i)-DPF_{proj}(i))^{2}}}{\sqrt{\frac{1}{N_{low}}\sum_{i=1}^{N_{low}}(DPF_{proj}(i))^{2}}}$ (24)
@@ 4.1 Result from experiment #1: setting of parameter $\alpha$
The results of Expe #1 are shown in Fig.2 On this figure, we report the distributions from the three measures detailed in Sec.3.1 ($Dev$, StdCrestStdCrest, and $Sep$) computed across the dataset #1, for different values of the $\alpha$ parameter.
Figure 2: Results from Expe #1. A: From top to bottom, we show the distributions of the three measures DevDev, StdCrestStdCrest, and SepSep across all the brains from the dataset #1 as violin plots, for the different values of $\alpha$. We highlight in gray the range of values for which the distributions are statistically equivalent to the one obtained using the best value for $\alpha$ (Wilcoxon-test with p-value>>.05). B: For each measure, we show the distributions obtained for the method SULC, the mean curvature and the classical DPFDPF from auzias_deep_2015, for comparison.
@@
From the plot of the normalized standard deviation on the crests $StdCrest$, we observe a strong decrease with increasing $\alpha$, from .24 for $\alpha=0$ to .12 for $\alpha=150$. The lowest median is obtained for $\alpha=400$, and the statistical tests show that the distributions are equivalent for $\alpha\in[200,1000]$.
From the plot of the difference between the median of the crest and the median of the fundi $Sep$, we observe a progressive increase with increasing values of $\alpha$. From a value of .2 for $\alpha=0$, the median of $Sep$ increases up to .9 for $\alpha=500$, with a highest median for $\alpha=2000$. The statistical tests indicate that the distributions are equivalent for $\alpha$ values above 1000.
@@
When comparing with the SULC and classical $DPF$ methods in Fig.1.B, we observe that their angular deviation $Dev$ are also quite low compared to mean curvature, and not different from the best setting of $\alpha$ for $DPF^{*}$. These distributions confirm the visual observation of the appropriate setting of the default parameters of SULC for brains having a global scale not too different from an adult. Note however that for $StdCrest$ and $Sep$, the distributions from $DPF^{*}$ with the best $\alpha$ are statistically different from SULC, and from the classical $DPF$.
Overall, we observe that the influence of $\alpha$ on the estimated suclal depth from $DPF^{*}$ is continuous, since a small variation of $\alpha$ will have a very small influence. In addition, $DPF^{*}$ shows better measures than both DPF and SULC for a large range of values of $\alpha$. Considering an optimal value $\alpha=50$ for $Dev$, $\alpha=400$ for $StdCrest$, and the plateau after $\alpha=1000$ for $Sep$, we choose an intermediate value of $\alpha=500$ for the rest of the experiments. This choice was corroborated by visual inspections.
@@ 4.2 Result from experiment #2: invariance to scaling
Figure 3: Results from Expe #2. A. The scaled surfaces obtained as described in Sec.3.2. B. Depth maps generated using the $DPF^{*}$ method (with $\alpha=500$). The scale-invariance is confirmed empirically: the estimated depth values are not influenced by the scaling. The correlation is thus perfect for all surfaces and the regression lines are super-imposed. C. Left: Depth estimations resulting from the SULC method applied to the scaled surfaces (y-axis) with respect to SULC on the initial mesh (x-axis). Both the slope and the correlation values depend on the scaling factor. Right: Spatial distribution of residuals from the linear regression models derived from the left plot. Consistently with the regression plot, the residuals increase with the scaling factor. The regions corresponding to most concave and convex geometric configurations show higher magnitude of error.
First, this experiment allows us to confirm empirically that the $DPF^{*}$ is scale-invariant: $D^{*}(M,s)=D^{*}(M,1)$ for s=2,3,4,5s={2,3,4,5}. This experimental confirmation aligns with the theoretical formulation we introduced earlier.
@@ 4.3 Result from experiment #3: Application to a large dataset
The results of Expe #3 are detailed in Fig.4 and corresponding caption. The distributions of the different sulcal depth estimation techniques (top row) are indicative of the influence of the global brain size. For SULC, on the left, the impact of the increase in brain size from left to right is visible on all the centiles. More specifically, the jump in the dispersion of the distributions (esp. visible on the 75th and 95th centiles) corresponds to the transition from the 879 newborns from the dHCP to the adults from the HCP. On the dHCP dataset, the dispersion progressively increases with brain size, while on the HCP dataset, the distributions are very similar across individuals despite variations in global brain size (52% of variation between the volumes of the smallest and largest adults from HCP).
@@
For the $DPF^{*}_{abs}$, on the right, we see that by multiplying the $DPF^{*}$ by the characteristic length as in Eq.14, we obtain an absolute estimation of the sulcal depth with a clear influence of the global size of the mesh, as expected. The increase in the dispersion of the distributions with brain size is clearly visible and progressive. Of note, the influence of brain size is stronger than for SULC on the adult subjects from the HCP dataset. The $DPF^{*}_{abs}$ shows a higher sensitivity than SULC to the 52% variation in brain volume across adults.
@@
Figure 4: Analysis of the distribution of SULC, $DPF^{*}$ and $DPF^{*}_{abs}$ across the 879 surfaces from the dHCP newborns and 1108 surfaces from HCP young adults. From left to right, the three columns correspond resp. to SULC, $DPF^{*}$ and $DPF^{*}_{abs}$. On the top row, we show the distribution of each sulcal depth estimation for all subjects, ordered following their global size. For each method, we show the mean in purple, the median in blue, the 25th and 75th centiles in green, and the 5th and 95th centiles in orange. On the second row, we show the corresponding matrices of the Wasserstein distance between every pair of surfaces. On the third row, we show for each method the distributions of the Wasserstein distances within the four sub-populations corresponding to the blocks illustrated on the matrix for the method SULC. At the bottom, we show for each method the KS statistic comparing the distribution in the block corresponding to the subgroup of largest brains, to the other blocks corresponding to smaller brains. See the text for further description.
4.4 Result from Experiment #4 : Quantifying brain folding dynamics
We show in Fig. 5 the 23 sulcal regions on the surfacic atlas from the dHCP bozek_construction_2018 at 28, 32 and 44 weeks post conception, for the left hemisphere. We provide in Appendix 6 the same illustration for the right hemisphere, as well as an illustration of the estimated suclal depth mapped onto the atlas. See de_vareilles_development_2023 for the complete description of the nomenclature used to label the sulci. In the bottom panel, we show the average of the sulcal depth within each sulcal region as a function of the age for SULC on left and $DPF_{abs}^{*}$ on the right. The color of each curve corresponds to the sulcus nomenclature shown at the top right of the figure. From the literature of early brain development (see de_vareilles_development_2023 for a review of the timing of formation of the sulci), we expect to observe 1) a decrease in the metrics reflecting the depth for all sulci and 2) a specific order across sulci in terms of dynamics of the deepening (some sulci becoming deeper before others) and in terms of relative overall depth (some sulci being deeper than others at all ages). The two observations are confirmed for $DPF_{abs}^{*}$ (on the right): according to de_vareilles_development_2023, the sulci colored in levels of blue are expected to appear before and be deeper than the sulci in red, which are expected to appear before the sulci in green. In addition, the decrease in depth metrics is clear for all sulci except for later forming sulci such as S.T.i and S.Rh for which the curve stays flat until 34-36 weeks and then the depth metric starts to decrease. The expected ordering of sulci according to their overall depth is well respected, with blue curves corresponding to deeper sulci systematically located below the red and green curves.
Regarding the SULC method (on the left), this ordering according to the overall depth is relatively preserved, but with noticeable exceptions such as the Olfactory Sulcus (S.Olf, shown in dark blue), for which the curve is located within the group of green curves (shallow folds) and above the red curves that are supposed to be shallower than S.Olf. In addition, we observe an unexpected reduction in depth with age (increase in corresponding curves) for the Calcarine Fissure (F.Cal, in cyan) and the Parieto-Occipital Fissure (F.P.O in light blue). These two sulci are expected to be already present at 28 weeks, but we expect they become deeper on this period, since the global brain size increases.
In this experiment, we thus observe that the $DPF_{abs}^{*}$ allows us to characterize the dynamics of the formation of main cortical sulci. The biological interpretation is more consistent with the literature than with SULC.
Figure 5: Top: Labeled sulcal regions shown on three time steps of the spatio-temporal atlas. The nomenclature of sulci and associated colors are shown on the right. The colors are indicative of the expected timing of formation: sulci colored in levels of blue are expected to appear before the sulci in red that appear before the sulci in green. Bottom: Curves showing the average sulcal depth estimated in the different sulcal regions for SULC (left) and $DPF_{abs}^{*}$ (Right).
4.5 Result from experiment #5 : Assessment of the potential influence of the spatial resolution of the mesh
We report in Table 1 the comparison between the sulcal depth estimated on the resampled surfaces and the estimation from the original mesh. We report the mean and standard deviation of the NRMSE across the 16 subjects. We observe that the influence of the resolution is limited, with differences reaching only 15 % for meshes that have been down-sampled by a factor 4.
Resolution 75 % 50 % 25 %
NRMSE ($\pm$ std) 3.1 $\pm$ 0.4 6.3 $\pm$ 0.6 15.4 $\pm$ 0.9
Table 1: Evolution of the mean and std of NRMSE with the spatial resolution
In addition, we note that the computation time of DPFDPF* is less than 1 minute for a mesh of  60k vertices on a regular laptop.